OpenStudy (anonymous):

Suppose that you take out an unsubsidized Stafford loan on September 1 before your junior year for $4500 and plan to begin paying it back on December 1 after graduation and grace period 27 months later. The interest rate is 6.8%. How much of what you will owe will be interest?

OpenStudy (anonymous):

i believe its $661.76

myininaya (myininaya):

cantorset can we use induction on your problem? do you know what that is?.

OpenStudy (anonymous):

yes we can

OpenStudy (anonymous):

how did you find me, lol

myininaya (myininaya):

i went to the homepage and you can see what everyone is doing

OpenStudy (anonymous):

the chat program doesnt work for me

OpenStudy (anonymous):

ok lets do this one for all n>=5 , n^3 < 3^n

myininaya (myininaya):

ok

myininaya (myininaya):

for n=5, we have 5^3=125<243=3^5. Sold the inequality holds for n=5

OpenStudy (anonymous):

right, basis case

myininaya (myininaya):

now we should assume it is true for integers k>=5 Then we show it is true for k+1 the induction step is the hardest so let me think about it

myininaya (myininaya):

it seems that you are going over logs so i'm pretty sure we may have to involved that somehow

OpenStudy (anonymous):

hmmm, they may be unrelated though. here is what i got ,

OpenStudy (anonymous):

that was a different question

OpenStudy (anonymous):

is there a way to post without having to click my mouse on post?

myininaya (myininaya):

no i guess not

OpenStudy (anonymous):

show that (k+1)^3 < 3^(k+1)

OpenStudy (anonymous):

k^3 + 3k^2 + 3k + 1 < 3^(k+1)

myininaya (myininaya):

omg i got it

myininaya (myininaya):

i got so excited i clicked out sorry. so we have (k+1)^3<K^3+1^3=k^3+1<3^k+1

OpenStudy (anonymous):

k ^3 + 3k^2 + 3k + 1 < k^3 + 3k^2 + 3k + 3k = k^3 + 3k^2 + 6k < k^3 + k^3 + k^3 = 3k^3 < 3*3^k = 3^(k+1)

OpenStudy (anonymous):

(k+1)^3 = k ^3 + 3k^2 + 3k + 1 < k^3 + 3k^2 + 3k + 3k = k^3 + 3k^2 + 6k < k^3 + k^3 + k^3 = 3k^3 < 3*3^k = 3^(k+1)

myininaya (myininaya):

i got excited ignore what i wrote

OpenStudy (anonymous):

(k+1)^3<K^3+1^3=k^3+1<3^k+1 this is false

OpenStudy (anonymous):

my proof isnt elegant, but i think it works, might have to do some side induction arguments for the intermediate steps

myininaya (myininaya):

you are right i got excited

OpenStudy (anonymous):

also you better use parenthesses :) more often

OpenStudy (anonymous):

the log proof is A LOT easier

myininaya (myininaya):

yep

OpenStudy (anonymous):

what did you get? log_3 n < n if n < 3^n , assuming we raise both sides to 3

OpenStudy (anonymous):

or you can start, if we can show that n < 3^n , for n>=1, then taking the log base 3 of both sides we have our result. what was your logic?

myininaya (myininaya):

ok lets skip the case for n=1 that is easy to do and we know it holds. so assume for integer k, log3(k)<k (3is base) now we want to show log3(k+1)<k+1 so we have ...

OpenStudy (anonymous):

so that proof goes, 1 < 3^n, assume k < 3^k then add 1 to both sides of that inequality we have k + 1 < 3^k + 1 < 3^k + 3 = 3^(k+1)

myininaya (myininaya):

yea maybe we should base 3 on both sides to get rid of log

OpenStudy (anonymous):

you mean exponentiate?

OpenStudy (anonymous):

you cant get rid of log with log :)

OpenStudy (anonymous):

can you help me with surface area of integral

myininaya (myininaya):

surface area of a solid?

OpenStudy (anonymous):

surface area of a torus

OpenStudy (anonymous):

so you cant finish the proof with log, no problemo

myininaya (myininaya):

one thing at a time

OpenStudy (anonymous):

how do you find the question i posted earlier

OpenStudy (anonymous):

this website is so confusing

myininaya (myininaya):

are you given the equation for the torus or is it a parametric curve/

OpenStudy (anonymous):

Calculus question: surface area of a torus. it is generated by circle with radius r , off center. Suppose the circle is centered at (x - R)^2 + y^2 = r^2, and R> r. so the distance from (0,0) to the center of the circle is (R,0) which has radius r. Ok the integral i got is int 2pi y sqrt [ 1+ (dy/dx)^2] = int 2pi sqrt ( r^2 - ( x-R)^2) [ 1 + (-2(x-R)/(2sqrt( r^2 - ( x-R)^2) ^2 ]

OpenStudy (anonymous):

we are given a circle , and its going to generate the torus by revolving about the y axis

OpenStudy (anonymous):

the final answer should be 4pi^2 R*r

myininaya (myininaya):

so that last thing you have the square cancels the sqrt

myininaya (myininaya):

ok im just going to do it on paper and scan it and you can open it give a few min

OpenStudy (anonymous):

right, good :) ok here is what i got

OpenStudy (anonymous):

integral 2pi y dS, remember ds can be sqrt 1 + dx/dy ^2 or sqrt 1 + dy/dx ^2

myininaya (myininaya):

we need another pi in there somewhere

myininaya (myininaya):

since are answer has pi^2

myininaya (myininaya):

this is what i got using the integral you found we can get times another pi it would be correct

myininaya (myininaya):

don't look at the name of the file. i know we are not finding volume

OpenStudy (anonymous):

right, its surface area, here lets use twiddla

OpenStudy (anonymous):

http://www.twiddla.com/521818

OpenStudy (anonymous):

if you dont mind, we can draw on whiteboard